OpenStudy (anonymous):

2/3(y+2) - 5/6(y-1) = 5/4(y-3) - 4 help?? I got 7, but not sure if its correct...

OpenStudy (amistre64):

2/3(y+2) - 5/6(y-1) = 5/4(y-3) - 4 2y/3 +4/3 -5y/6 +5/6 = 5y/4 -15/4 -4 ; start clearing fractions

OpenStudy (amistre64):

2y/3 +4/3 -5y/6 +5/6 = 5y/4 -15/4 -4 ; *3 2y +4 -15y/6 +5/6 = 15y/4 - 45/4 -12 ; *6 12y +24 -15y +5 = 90y/4 -270/4 -72

OpenStudy (amistre64):

12y +24 -15y +5 = 90y/4 -270/4 -72 ; *4 48y +96 -60y +20 = 90y -270 -288 ; combine like terms

OpenStudy (amistre64):

48y -60y +20+96 = 90y -558 -12y -90y = -558 -116 -102y = -674 y = -674/-102 = 6.607.... if i kept track of it all

OpenStudy (anonymous):

Oh wow! lol! Well, I came up with the wrong answer then :/ thanks for taking the time to explain!! :)

OpenStudy (amistre64):

y = 7 is correct; so I miscalced someplace :)

OpenStudy (anonymous):

Ohh lol! :) Thank youu

OpenStudy (amistre64):

2y +4 -15y/6 +5/6 <-- right there :) 5/6 * 3 doesnt equal 5/6 :)

OpenStudy (anonymous):

Ohhh... Gotcha ;)

OpenStudy (amistre64):

lets see if I can get it back on track :) 48y -60y +60+96 = 90y -558 -12y -90y = -558 -156 -102y = -714 y = -714/-102 = 7 .... yay!!

OpenStudy (anonymous):

Woo Hoo!! :) haha thank you! ur a lifesaver